# RS-MoE: Mixture of Experts for Remote Sensing Image Captioning and Visual Question Answering

### v1

@@ III-C Model Tuning
As illustrated in Figure 1, the tuning process of the proposed RS-MoE model is divided into two stages. Considering the complexity and diversity of information within remote sensing images, directly training our model can lead to performance degradation due to sparsity. This sparsity stems from the inadequate representation of remote sensing-specific features in models pre-trained on natural images, creating inefficiencies in capturing the intricate characteristics of remote sensing imagery, ultimately impacting model performance. To address this, we propose a two-stage training strategy. The first stage begins by independently training the VLM Encoder and LLM, ensuring that the encoder can extract rich and detailed visual features specific to remote sensing imagery while enhancing its capability to generate meaningful, context-aware feature representations. Once the VLM Encoder generates stable visual features, the second stage involves fine-tuning the language models in the MoE. In this stage, focus shifts to adjusting the MoE Block’s weights to produce more detailed captions tailored specifically for RSIC tasks, with each language model concentrating on its designated task.
@@ III-C1 Stage I
The LLM block receives output features and instructions from the instruction-aware VLM Encoder to generate the final captions for remote sensing images, which plays a crucial role in synthesizing the contextual data provided by the VLM Encoder into coherent and detailed textual descriptions. The fine-tuning process is focused on enhancing the model’s ability to understand the complex visual patterns typical of remote sensing images and to translate them into accurate linguistic representations. By tuning the LLM Block, the model becomes more efficient at generating captions that are deeply aligned with the specific requirements of remote sensing tasks, including adapting to various landscapes, recognizing subtle environmental changes, and describing intricate relationships between different geographical features in the images. Similarly, we employ the LoRA technique to reduce the number of trainable parameters, thereby enhancing the model’s efficiency and manageability during the training process. LoRA is specifically applied to the query and value layers within the attention module, as well as to the feed-forward networks. This optimization not only streamlines the training process but also sharpens the focus of the attention mechanisms, improving the overall processing and responsiveness of the model.
@@ III-C2 Stage II
In this stage, our focus shifts to fine-tuning the MoE Block’s weights to produce more detailed captions specifically for remote sensing image captioning tasks. Tuning the MoE Block involves detailed adjustments to the interaction mechanisms between the distinct LLMs to ensure that each LLM can effectively respond to its specific prompt by extracting and interpreting nuanced visual cues. This fine-tuning facilitates the development of a deep, segmented understanding of the image, allowing each expert within the MoE to generate a part of the caption that reflects its specialized knowledge of the image’s theme, objects, or relationships. As these parts come together, they form a comprehensive and detailed caption that captures both the breadth and depth of the remote sensing imagery.
@@ IV-A1 RSICap
This dataset is built on the DOTA object detection dataset, which contains 2,585 high-quality remote sensing image-text pairs for training vision-language models on remote sensing images. The images are randomly chosen from the DOTA dataset and resized to 512 $\times$ 512 pixels. The text annotations are manually labeled by five experts in the remote sensing field. ç To evaluate the VLM model on this dataset, the authors further constructed an evaluation set, RSIEval, that consists of 1,000 512x512 remote sensing images with manual annotations.
UCM-Captions
@@
RSICD
This dataset contains 10,921 remote sensing images collected from four platforms: Google Earth, Baidu Maps, MapABC, and Tianditu. The images are uniformly resized to 224 $\times$ 224 pixels, although they come in various resolutions. The captions for these images were created by volunteers who possess experience in annotation and knowledge of remote sensing. For each image, the original caption varies from one to five sentences. To enrich the descriptions, the authors duplicated the captions that contained fewer than five sentences to ensure that each image had five distinct sentences in its caption.
@@ IV-D Results
IV-D1 Results on RSICap
IV-D2 Results on UCM-Captions
To further illustrate the effectiveness of our model in generating accurate and contextually aware captions, we selected two remote sensing images with distinct scenes from the test set of the UCM-Captions dataset. The captions generated by RSGPT and our model for these images are shown in Figure 3. As observed, both models successfully capture the overall themes of the images. However, for detailed descriptions, particularly in the first remote sensing image, the captions generated by RSGPT exhibit discrepancies with the actual scene. In contrast, the captions produced by the proposed RS-MoE model provide precise and detailed representations of both the objects and their spatial relationships within the remote sensing images, underscoring our model’s advantage in understanding complex spatial arrangements in diverse remote sensing imagery.
@@ IV-D2 Results on UCM-Captions
IV-D3 Results on Sydney-Captions
To further assess our model’s efficacy for the RSIC task, we extended the evaluation to the traditional Sydney-Captions dataset. In this experiment, sixteen deep learning models specifically designed for remote sensing image captioning, along with two recent VLM models, RSGPT [15] and RS-CapRet [41], were selected as comparative benchmarks. Noted that we still employed the 7B variant of RS-MoE without any additional fine-tuning on the Sydney-Captions dataset. The comparative performance of these baseline models and our model is detailed in Table III. Analysis of the results demonstrates that, even without fine-tuning, our model consistently outperformed all competing models across all seven evaluation metrics. A particularly notable achievement is its performance on the METEOR metric, where it achieved a score of 56.23, surpassing the next best model, which scored 41.37, by a substantial margin of 15%.
To further illustrate the captioning capabilities of our model in comparison to RSGPT, we randomly selected two remote sensing images from the test set of the Sydney-Captions dataset, as shown in Figure 4. These images depict scenes of a residential area and an airport. Figure 4 shows that while both RSGPT and RS-MoE provide generally accurate descriptions of the overall themes and specific details, the RSGPT-generated captions lack the depth observed in our model’s outputs. Additionally, the last sentence generated by RSGPT for the first image remains incomplete, underscoring our model’s advantage in generating coherent and contextually rich descriptions.
@@ IV-D3 Results on Sydney-Captions
TABLE V: Comparison results of different numbers of LLMs in the MoE block on the RSICap dataset for the RSIC task.
@@
TABLE VI: Comparison results of different LLM model in the MoE block on the RSICap dataset for the RSIC task.
@@
IV-D4 Results on RSICD
To further illustrate the effectiveness of the generated captions, we selected two remote sensing images from the test set of the RSICD dataset and displayed the captioning results generated by RSGPT and our model in Figure 5. It is evident from Figure 5 that, unlike RSGPT, our RS-MoE model accurately describes the spatial positions of objects within the images. This capability demonstrates RS-MoE’s enhanced ability to capture and articulate spatial relationships and arrangements, which are essential for a comprehensive understanding of remote sensing imagery. By providing detailed, contextually rich descriptions that reflect the actual scene, RS-MoE ensures that the generated captions are both informative and precise, thereby improving the overall interpretability of remote sensing data.
V Discussion
In this section, we conduct a comprehensive analysis of the effectiveness and generalizability of the proposed RS-MoE model from two perspectives. Firstly, we verify the effectiveness of our model and training strategy through three specific ablation studies. Specifically, the first ablation study investigates the impact of varying the number of LLMs within the MoE block. The second study assesses the effects of using different types of LLMs on model performance. The third ablation study focuses on evaluating the effectiveness of our two-stage training strategy. Subsequently, we extend the application of our model to the Remote Sensing Visual Question Answering (RSVQA) task to assess its broader applicability and generalization capabilities across diverse contexts.
TABLE VII: Comparison results of different training strategy for our model on the RSICap dataset for the RSIC task.
@@ V Discussion
V-A Ablation Study
V-A1 Effect of Different Numbers of LLMs in the MoE Block
@@
We evaluated the performance of these models on the RSIC task using the RSIEval dataset, with detailed results presented in Table V. By comparing the results for configurations with more than one LLM to those with only a single LLM, we observe substantial performance improvements when the task is divided into subtasks. This results further validates the effectiveness of incorporating the idea of MoE into the remote sensing image captioning task. Additionally, comparing the results for models with 2, 3, and 4 LLMs, we found that using three LLMs yields the best performance. Consequently, in our experiments, we employ three LLMs to focus on generating captions for the image’s theme, included objects, and spatial relationships among objects, highlighting the advantage of the MoE framework for handling complex, multi-faceted tasks in remote sensing. This setup not only enhances the interpretability of the generated captions but also demonstrates the flexibility and scalability of our model when addressing intricate details in remote sensing imagery.
TABLE VIII: Results of our model and state-of-the-art models on the RSVQA test set of RSIEval.
@@
V-A2 Effect of Different LLM Models
@@
Each of the seven RS-MoEmodel variants was trained as a baseline on the RSICap dataset and evaluated on the RSIEval dataset. The performance metrics are displayed in Table VI. As observed from the results, all selected lightweight LLMs under our MoE framework achieved reasonably satisfactory performance in remote sensing image captioning. This consistency across different LLMs demonstrates the robustness and adaptability of our proposed framework, indicating that even with various lightweight LLM configurations, the RS-MoEmodel is able to maintain effective and contextually relevant caption generation for complex remote sensing imagery. This suggests that the our MoE framework can be flexibly integrated with different LLMs, providing a versatile solution adaptable to the needs of various tasks and computational environments.
V-A3 Effect of Two-stage Training Strategy
@@
We further evaluate the effectiveness of our proposed two-stage training strategy. As explained in Section III-C, directly training our RS-MoEmodel can lead to performance degradation due to sparsity issues stemming from the distinct characteristics of remote sensing images compared to natural images. To validate this hypothesis, we implemented a baseline model trained using a single-stage approach, directly optimizing the model without employing the two-stage strategy. Comparative results for both training methods are shown in Table VII. The findings indicate that the two-stage training strategy significantly outperforms the single-stage approach, underscoring the advantages of this method. By incorporating the first training stage, the VLM encoder and LLM Block learn feature representations tailored to remote sensing imagery, gaining a foundational understanding of remote sensing-specific visual patterns and linguistic cues. This process establishes a robust basis for accurate feature extraction and preliminary captioning. This stage essentially fine-tunes the model’s ability to interpret complex spatial arrangements and thematic elements unique to remote sensing imagery, elements often underrepresented in models pretrained on natural images. In the second stage, the model builds upon this solid foundation, requiring only minimal fine-tuning to adapt the MoE block. Consequently, each LLM in the MoE block can effectively focus on its designated task, whether it involves capturing thematic elements, identifying specific geographic objects, or interpreting spatial relationships among objects. This task-oriented fine-tuning results in captions that are not only precise but also provide a comprehensive view of the remote sensing scene, with detailed and contextually relevant descriptions. Overall, this two-stage training strategy represents a core innovation of our work, demonstrating how carefully designed training frameworks can achieve superior results in the specialized domain of remote sensing.

### v2

@@ III-C Model Tuning
As illustrated in Fig. 1, the tuning process of the proposed RS-MoE model is divided into two stages. Considering the complexity and diversity of information within remote sensing images, directly training our model can lead to performance degradation due to sparsity. This sparsity stems from the inadequate representation of remote sensing-specific features in models pre-trained on natural images, creating inefficiencies in capturing the intricate characteristics of remote sensing imagery, ultimately impacting model performance. To address this, we propose a two-stage training strategy. The first stage begins by independently training the VLM Encoder and LLM, ensuring that the encoder can extract rich and detailed visual features specific to remote sensing imagery while enhancing its capability to generate meaningful, context-aware feature representations. Once the VLM Encoder generates stable visual features, the second stage involves fine-tuning the language models in the MoE. In this stage, focus shifts to adjusting the MoE Block’s weights to produce more detailed captions tailored specifically for RSIC tasks, with each language model concentrating on its designated task.
@@ III-C1 Stage I
The LLM block receives output features from the VLM Encoder and input instructions to generate the final captions for remote sensing images, which plays a crucial role in synthesizing the contextual data provided by the VLM Encoder into coherent and detailed textual descriptions. The fine-tuning process is focused on enhancing the model’s ability to understand the complex visual patterns typical of remote sensing images and to translate them into accurate linguistic representations. By tuning the LLM Block, the model becomes more efficient at generating captions that are deeply aligned with the specific requirements of remote sensing tasks, including adapting to various landscapes, recognizing subtle environmental changes, and describing intricate relationships between different geographical features in the images. Similarly, we employ the LoRA technique to reduce the number of trainable parameters, thereby enhancing the model’s efficiency and manageability during the training process. LoRA is specifically applied to the query and value layers within the attention module, as well as to the feed-forward networks. This optimization not only streamlines the training process but also sharpens the focus of the attention mechanisms, improving the overall processing and responsiveness of the model.
@@ III-C2 Stage II
In this stage, our focus shifts to fine-tuning the weights of the LLMs within the MoE Block to produce more detailed captions specifically for remote sensing image captioning tasks. Tuning the MoE Block involves detailed adjustments to the interaction mechanisms between the distinct LLMs to ensure that each LLM can effectively respond to its specific prompt by extracting and interpreting nuanced visual cues. This fine-tuning facilitates the development of a deep, segmented understanding of the image, allowing each expert within the MoE to generate a part of the caption that reflects its specialized knowledge of the image’s theme, objects, or relationships. As these parts come together, they form a comprehensive and detailed caption that captures both the breadth and depth of the remote sensing imagery.
@@ IV-A1 RSICap
This dataset is built on the DOTA object detection dataset, which contains 2,585 high-quality remote sensing image-text pairs for training vision-language models on remote sensing images. The images are randomly chosen from the DOTA dataset and resized to 512 $\times$ 512 pixels. The text annotations are manually labeled by five experts in the remote sensing field. To evaluate the VLM model on this dataset, the authors further constructed an evaluation set, RSIEval, that consists of 1,000 512x512 remote sensing images with manual annotations.
IV-A2 UCM-Captions
@@ IV-A3 Sydney-Captions
IV-A4 RSICD
This dataset contains 10,921 remote sensing images collected from four platforms: Google Earth, Baidu Maps, MapABC, and Tianditu. The images are uniformly resized to 224 $\times$ 224 pixels, although they come in various resolutions. The captions for these images were created by volunteers who possess experience in annotation and knowledge of remote sensing. For each image, the original caption varies from one to five sentences. The authors duplicated the captions that contained fewer than five sentences to ensure that each image had five distinct sentences in its caption.
IV-A5 VRSBench
This dataset consists of 29,614 images, each paired with a human-verified, detailed caption. These captions include abstract image attributes, such as source, resolution, and scene type, alongside object-specific details like quantity, color, shape, and relative positioning. They emphasize clear, observable features, excluding ambiguous elements, and may describe additional objects, such as buildings, roads, and trees, beyond those in the source datasets. Each caption typically comprises 3-7 sentences, averaging 54 words.
@@ IV-D Results
TABLE V: Results of our model and state-of-the-art models on the VRSBench dataset for RSIC Task.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE_L CIDEr CHAIR
GeoChat w/o ft [21] 13.9 6.6 3.0 1.4 7.8 13.2 0.4 0.42
GPT-4V [58] 37.2 22.5 13.7 8.6 20.9 30.1 19.1 0.83
MiniGPT-v2 [59] 36.8 22.4 13.9 8.7 17.1 30.8 21.4 0.73
LLaVA-1.5 [60] 48.1 31.5 21.2 14.7 21.9 36.9 33.9 0.78
GeoChat [21] 46.7 30.2 20.1 13.8 21.1 35.2 28.2 0.77
Mini-Gemini [61] 47.6 31.1 20.9 14.3 21.5 36.8 33.5 0.77
RS-MoE 55.3 34.9 25.5 17.2 25.1 42.8 40.5 0.81
IV-D1 Results on RSICap
IV-D2 Results on UCM-Captions
To further illustrate the effectiveness of our model in generating accurate and contextually aware captions, we selected two remote sensing images with distinct scenes from the test set of the UCM-Captions dataset. The captions generated by RSGPT and our model for these images are shown in Figure 3. As observed, both models successfully capture the overall themes of the images. However, for detailed descriptions, particularly in the first remote sensing image, the captions generated by RSGPT exhibit discrepancies with the actual scene. In contrast, the captions produced by RS-MoE model provide precise and detailed representations of both the objects and their spatial relationships within the remote sensing images, underscoring our model’s advantage in understanding complex spatial arrangements in diverse remote sensing imagery.
IV-D3 Results on Sydney-Captions
To further assess our model’s efficacy for the RSIC task, we extended the evaluation to the traditional Sydney-Captions dataset. In this experiment, sixteen deep learning models specifically designed for remote sensing image captioning, along with two recent VLM models, RSGPT [20] and RS-CapRet [56], were selected as comparative benchmarks. Noted that we still employed the 7B variant of RS-MoE without any additional fine-tuning on the Sydney-Captions dataset. The comparative performance of these baseline models and our model is detailed in Table III. Analysis of the results demonstrates that, even without fine-tuning, our model consistently outperformed all competing models across all seven evaluation metrics.A particularly notable achievement is its performance on the METEOR metric, where it achieved a score of 56.23, surpassing the next best model’s score of 41.37 by a significant margin of 15%.
To further illustrate the captioning capabilities of our model in comparison to RSGPT, we randomly selected two remote sensing images from the test set of the Sydney-Captions dataset, as shown in Figure 4. These images depict scenes of a residential area and an airport. Figure 4 shows that while both RSGPT and RS-MoE provide generally accurate descriptions of the overall themes and specific details, the RSGPT-generated captions lack the depth observed in our model’s outputs. Additionally, the last sentence generated by RSGPT for the first image remains incomplete, underscoring our model’s advantage in generating contextually descriptions.
TABLE VI: Comparison results of RS-MoE with and without the instruction router on the RSICap dataset for the RSIC task.
Training Strategy BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE_L CIDEr
RS-MoE w/o Instruction Router 47.32 33.18 25.66 21.42 22.87 25.33 102.51
RS-MoE w/ Instruction Router 82.13 65.44 51.93 42.55 40.28 35.72 158.36
TABLE VII: Comparison results of different numbers of LLMs in the MoE block on the RSICap dataset for the RSIC task.
@@ IV-D3 Results on Sydney-Captions
IV-D4 Results on RSICD
To further illustrate the effectiveness of the generated captions, we selected two remote sensing images from the test set of the RSICD dataset and displayed the captioning results generated by RSGPT and our model in Figure 5. It is evident from Figure 5 that, unlike RSGPT, our RS-MoE model accurately describes the spatial positions of objects within the images. This capability demonstrates RS-MoE’s enhanced ability to capture and articulate spatial relationships and arrangements, which are essential for a comprehensive understanding of remote sensing imagery. By providing detailed, contextually rich descriptions that reflect the actual scene, RS-MoE ensures that the generated captions are both informative and precise, thereby improving the overall interpretability.
IV-D5 Results on VRSBench
We further evaluate our method on the RSBench dataset, a benchmark specifically designed to test detailed captioning capabilities. In our experiments, we compare our method against several strong baselines, including GeoChat without finetuning [21], GPT-4V [58], MiniGPT-v2 [59], LLaVA-1.5 [60], GeoChat [21], and Mini-Gemini [61]. The results of these models are summarized in Table V. As shown in the table, our method demonstrates outstanding performance across all major metrics. Notably, our method achieves a BLEU-1 score of 55.3, significantly surpassing Mini-Gemini’s BLEU-1 score of 47.6. Similarly, our method achieves a BLEU-3 score of 25.5, outperforming Mini-Gemini’s 20.9. Although GPT-4V achieves the highest CHAIR score of 0.83, our method achieves a competitive score of 0.81, substantially outperforming other baselines. These results clearly demonstrate the effectiveness of our method in generating detailed and descriptive captions that capture both abstract image attributes and object-specific details. The substantial improvements over state-of-the-art baselines on RSBench further highlight the robustness and adaptability of our method in addressing the unique challenges of vision-language modeling for detailed captioning tasks.
V Discussion
In this section, we conduct a comprehensive analysis of the effectiveness and generalizability of the proposed RS-MoE model from two perspectives. Firstly, we verify the effectiveness of our model and training strategy through four ablation studies. These include examining the effect of the Instruction Router, varying the number of LLMs in the MoE block, using different types of LLMs, and evaluating the two-stage training strategy. Subsequently, we extend the application of our model to the Remote Sensing Visual Question Answering (RSVQA) task, testing its performance on two datasets, RSEval and RSIVQA, to assess its broader applicability and generalization capabilities.
V-A Ablation Study
V-A1 Effect of Instruction Router
To evaluate the effectiveness of the instruction router within the MoE Block, we conducted an ablation study comparing the model’s performance with and without the instruction router. For a fair comparison, both setups use three LLMs as expert models. When the instruction router is absent, the input instructions are directly passed to the LLMs without any dynamic prompt generation, and each LLM independently generates outputs.The results of the two models are shown in Table VI. It can be found that the model without the instruction router achieves a BLEU-1 score of 47.32, a METEOR score of 22.87, and a CIDEr score of 102.51. In contrast, incorporating the instruction router leads to significant improvements, with the model achieving a BLEU-1 score of 82.13, a METEOR score of 40.28, and a CIDEr score of 158.36. These results highlight the critical role of the instruction router in dynamically generating tailored prompts based on input instructions and visual features, effectively guiding the expert models to produce more precise and contextually appropriate descriptions. The substantial performance gains observed with the instruction router emphasize its importance in enabling the model to adapt to diverse input images and tasks. By leveraging Prompt Learning principles, the instruction router enhances the model’s ability to extract meaningful visual information and generate accurate, detailed descriptions, demonstrating its indispensability in the proposed framework.
TABLE VIII: Comparison results of different LLM model in the MoE block on the RSICap dataset for the RSIC task.
@@ V-A1 Effect of Instruction Router
TABLE IX: Comparison results of different training strategy for our model on the RSICap dataset for the RSIC task.
@@
V-A2 Effect of Different Numbers of LLMs in the MoE Block
@@
We evaluated the performance of these models on the RSIC task using the RSIEval dataset, with detailed results presented in Table VII. By comparing the results for configurations with more than one LLM to those with only a single LLM, we observe substantial performance improvements when the task is divided into subtasks. These results further validate the effectiveness of incorporating the idea of MoE into the remote sensing image captioning task. Additionally, comparing the results for models with 2, 3, and 4 LLMs, we found that using three LLMs yields the best performance. Consequently, in our experiments, we employ three LLMs to focus on generating captions for the image’s theme, included objects, and spatial relationships among objects, highlighting the advantage of the MoE framework for handling complex, multi-faceted tasks in remote sensing. This setup not only enhances the interpretability of the generated captions but also demonstrates the flexibility and scalability of our model when addressing intricate details in remote sensing imagery.
TABLE X: Results of our model and state-of-the-art models on the RSVQA test set of RSIEval for RSVQA task.
@@
V-A3 Effect of Different LLM Models
@@
Each of the seven RS-MoEmodel variants was trained as a baseline on the RSICap dataset and evaluated on the RSIEval dataset. The performance metrics are displayed in Table VIII. As observed from the results, all selected lightweight LLMs under our MoE framework achieved reasonably satisfactory performance in remote sensing image captioning. This consistency across different LLMs demonstrates the robustness and adaptability of our proposed framework, indicating that even with various lightweight LLM configurations, the RS-MoEmodel is able to maintain effective and contextually relevant caption generation for complex remote sensing imagery. This suggests that our MoE framework can be flexibly integrated with different LLMs, providing a versatile solution adaptable to the needs of various tasks and computational environments.
V-A4 Effect of Two-stage Training Strategy
@@
We further evaluate the effectiveness of our proposed two-stage training strategy. As explained in Section III-C, directly training our RS-MoEmodel can lead to performance degradation due to sparsity issues stemming from the distinct characteristics of remote sensing images compared to natural images. To validate this hypothesis, we implemented a baseline model trained using a single-stage approach, directly optimizing the model without employing the two-stage strategy. Comparative results for both training methods are shown in Table IX. The findings indicate that the two-stage training strategy significantly outperforms the single-stage approach, underscoring the advantages of this method. By incorporating the first training stage, the VLM encoder and LLM Block learn feature representations tailored to remote sensing imagery, gaining a foundational understanding of remote sensing-specific visual patterns and linguistic cues. This process establishes a robust basis for accurate feature extraction and preliminary captioning. This stage essentially fine-tunes the model’s ability to interpret complex spatial arrangements and thematic elements unique to remote sensing imagery, elements often underrepresented in models pretrained on natural images. In the second stage, the model builds upon this solid foundation, requiring only minimal fine-tuning to adapt the MoE block. Consequently, each LLM in the MoE block can effectively focus on its designated task, whether it involves capturing thematic elements, identifying specific geographic objects, or interpreting spatial relationships among objects. This task-oriented fine-tuning results in captions that are not only precise but also provide a comprehensive view of the remote sensing scene, with detailed and contextually relevant descriptions. Overall, this two-stage training strategy represents a core innovation of our work, demonstrating how carefully designed training frameworks can achieve superior results in the specialized domain of remote sensing.
TABLE XI: Results of our model and state-of-the-art models on the RSIVQA dataset for RSVQA task.
Model Yes/No Number Others Average Accuracy Overall Accuracy
MAIN [66] 92.82 56.71 54.50 68.01 77.39
EasyToHard [67] 95.49 49.03 63.65 69.39 79.70
SHRNet [68] 97.64 57.89 84.60 80.04 84.46
RSAdapter [69] 97.90 62.64 92.47 84.34 87.10
RS-MoE 98.01 75.46 92.88 88.78 89.12